OpenStudy (anonymous):

PRE-CAL; use the given pair of functions to find and simplify expressions for the following functions and state the domain of each using interval notation. • (gf)(x) • (f g)(x) • (f f)(x) 4. f(x)= 2x/x^2-4 and g(x)=√1-

OpenStudy (anonymous):

I don't understand the number 4 at the last sentence represent for what? 4f(x) ?

OpenStudy (anonymous):

@Hoa i apologize that was the question number

OpenStudy (anonymous):

for gf(x), at g function, wherever you see x, replace that x by f(x) and then solve as usual. the same with other parts. you may confuse at ff(x). however, nothing to be confused , do as usual, f (fx) = 2(fx) / f(x) ^2 -4 =\[ \frac{ 2\frac{ 2x }{ x^2-4 } }{(\frac{ 2x }{ x^2-4 })^2 -4 }\]

OpenStudy (anonymous):

do step by step you can solve it. hope this helps